If $7500 is invested at 6% annual simple interest, how much should be invested at 9% annual simple interest so that the total yearly income from both investments is $882.
Answer by lwsshak3(11628)

If $7500 is invested at 6% annual simple interest, how much should be invested at 9% annual simple interest so that the total yearly income from both investments is $882.
**
let x=$ invested at 9%
..
6%*7500+9%x=882
450+.09x=882
.09x=882-450=432
x=432/.09=4800
Ans:
$4800 should be invested at 9%